Find the probability that Х falls in the shaded area. Х 0 1 2 3 4 5 6 7 8 9 10
Answer #1

X 5 height Yg base of a Rechangle = (5-8)= 5 Area of a Rechangle base & height (5-0) G) 5 x 0.625 ploexas) = 0.625